# Evaluating the Performance of ChatGPT for Spam Email Detection

### v1

@@ Abstract
Email continues to be a pivotal and extensively utilized communication medium within professional and commercial domains. Nonetheless, the prevalence of spam emails poses a significant challenge for users, disrupting their daily routines and diminishing productivity. Consequently, accurately identifying and filtering spam based on content has become crucial for cybersecurity. Recent advancements in natural language processing, particularly with large language models like ChatGPT, have shown remarkable performance in tasks such as question answering and text generation. However, its potential in spam identification remains underexplored. To fill in the gap, this study attempts to evaluate ChatGPT’s capabilities for spam identification in both English and Chinese email datasets. We employ ChatGPT for spam email detection using in-context learning, which requires a prompt instruction and a few demonstrations. We also investigate how the training example size affects the performance of ChatGPT. For comparison, we also implement five popular benchmark methods, including naive Bayes, support vector machines (SVM), logistic regression (LR), feedforward dense neural networks (DNN), and BERT classifiers. Though extensive experiments, the performance of ChatGPT is significantly worse than deep supervised learning methods in the large English dataset, while it presents superior performance on the low-resourced Chinese dataset, even outperforming BERT in this case.
Index Terms:
@@
ChatGPT, Large Language Models, Spam email detection, Cybersecurity, In-context Learning
I Introduction
@@
Used as a swift and cost-efficient communication tool, email underpins global connectivity and is fundamental in various professional and personal activities, such as international electronic commerce [1]. As of 2020, the international email user base was approximately 4.037 billion. Due to its very nature of accessibility, email inevitably compromises on privacy [2], hence, exploited by corporations, organizations, and criminals alike for marketing and fraudulent purposes.
@@
Spam content predominantly encompasses deceptive information or advertisements [3], and occasionally includes harmful elements such as computer viruses, explicit material, or adverse information. The rampant spread of spam not only poses a nuisance to individuals but also instigates a squandering of network resources, in addition to eliciting a myriad of security threats [4]. As such, the development of an efficacious model capable of accurately identifying and consequently filtering spam is of paramount importance [5].
@@
Numerous methods have been established in the realm of spam detection and classification [6; 7]. Conventional approaches encompass techniques such as character matching, the naive Bayes algorithm [8], support vector machines (SVM) [9] and ensemble learning [10]. In the wake of deep learning advancements, a variety of deep models, including Convolutional Neural Networks (CNN), Recurrent Neural Networks (RNN), and BERT, have been employed in spam recognition [11; 12; 13]. For instance, [14] proposed a comprehensive dual-layer architecture for phishing and spam email detection. [15] developed a hierarchical attention hybrid deep learning method for spam email detection.
@@
Recently large language models (LLMs) such as ChatGPT have shown impressive performance in many natural language processing tasks [16], including question answering, text generation, etc. In-context learning [17; 18], central to these large models, refers to their prediction capabilities given a specific context; these models analyze and generate words or sentences based on the instruction (i.e., prompt), which aids their ability to coherently operate in tasks such as conversation, text completion, translation, and summarization. However, while ChatGPT has been used in various tasks of text processing, little research has been done on spam detection. To fill in the gap, we aim to explore how well ChatGPT performs at spam detection via in-context learning. We design experiments to systematically evaluate the performance of ChatGPT in both English and Chinese email datasets.
@@
II Related Works
II-A Spam Classification
@@
Based on the types of data used, existing spam classification algorithms can be broadly categorized into two groups: The first category involves recognition based on mail features. This approach involves extracting features from mail titles and employing supervised classifiers like the SVM model for classification [19]. Some studies have also considered the cumulative relative frequency of each spam feature to enhance the performance of supervised classifiers [12]. The feature-based spam recognition dataset is multidimensional, making it suitable for multiple classification models. Consequently, researchers have compared the performance of different machine learning models [20], such as SVM, deep neural networks (DNN), and k-nearest neighbor (KNN) [21], in spam recognition tasks.
@@
The second category focuses on email content identification. For instance, the relaxed online SVM model can be utilized to effectively filter and classify mail content [22], naive Bayes classifiers have also demonstrated effectiveness in email classification [23]. In recent years, deep learning algorithms have been increasingly employed in spam recognition. CNNs [11], for example, have proven to be effective in identifying spam. The transfer learning of BERT models has also been employed for spam classification, yielding high accuracy [12]. Long Short-Term Memory (LSTM) models have shown better performance on short messages (SMS) datasets compared to machine learning classification algorithms like random forest and SVM [24].
II-B GPT and Its Applications
@@
GPT [25] is one of the landmark models proposed on the basis of Transformer proposed by Vaswani in 2017 [26], another representative model is the BERT model proposed by [27]. On this basis, GPT2 [16] and GPT3 [28] have been proposed and widely used. Excellent performance was achieved tasks such as text generation, question and answer, translation, summary, and dialogue.
@@
ChatGPT [29] is an upgraded version of OpenAI in 2022 based on the GPT-3.5 model, which is optimized for conversation tasks, adding input and output of conversation history, and control of conversation policies. ChatGPT excels at conversational tasks and can have natural and fluid conversations with humans.
@@
Spam recognition is a text classification task. ChatGPT generally performs well on text classification tasks, and researchers have demonstrated that the model performs well on some text classification tasks [30]. GPT can be used to assist classification decision, by using chatgpt to extract structured and refined knowledge from knowledge graph as the characteristics of classification tasks. [31]. It can also be used directly for text classification with appropriate Prompt design, such as for agricultural [32], business news [33] and even medical dialogues [34].
@@
A few research have explored the application of ChatGPT in text binary classification [34], which can provide some reference for spam recognition. In recent studies, ChatGPT was able to identify opinions, sentiment, and emotions contained in the text and categorize emotions. In addition, GPT also performed well in sentiment analysis of disease-related papers [35] and Twitter texts [36], indicating that ChatGPT is capable of a wide range of sentiment analysis tasks for texts in different fields.
@@
Due to some uncertainty in the answers of large language models, [37], when using chatgpt for text classification tasks, it is necessary to carefully design hints and related strategies to make the answers as stable and effective as possible. If necessary, a small number of samples can be added to guide the model [38], such as One-shot Prompting [39],Few-shot Prompting [40], and Chain-of-Thought Prompting(CoT) design [41].
III Experimental Setup
@@ III-A Datasets
TABLE I: Detail of Datasets
@@
The general information of the dataset is shown in Table I. This table displays the size of two datasets, and the number of spam and ham emails in these datasets.
III-B In-context Learning
To utilize ChatGPT via in-context learning, it is necessary to design an instruction, the so-called prompt. We have tested two prompts in this research.
III-B1 Base Zero-shot Prompt
@@
Forget all your previous instructions, pretend you are a junk mail detection expert who try to identify whether a mail is spam mail. Forget all your previous instructions, pretend you are a junk mail detection expert who try to identify whether a mail is spam mail. Answer “spam” if it is spam, “ham” if not. Now you can identify whether the mail is spam:
@@
III-B2 Enhanced Zero-shot Prompt
Chatgpt’s performance is sometimes inconsistent, and we can enhance the classification efficiency of the model by adjusting the design of prompt [42]. On top of the observation from real-world questions, we propose an enhanced prompt template:
@@ III-B2 Enhanced Zero-shot Prompt
We made this adjustment because ChatGPT sometimes had difficulty identifying email categories and gave responses like ”neutral” or ”ambiguous” during training on real questions. Therefore, in order to achieve better binary classification, enhanced prompt produces clearer classification results.
III-C Baselines
@@
  • Naive Bayes: It is a classification algorithm based on Bayes’ theorem and feature independence assumption [23]. It assumes that input features are independent of each other and uses prior probabilities of classes and conditional probabilities of features to calculate posterior probabilities for classification. Naive Bayes classifier is simple, computationally efficient, and widely used in text classification and spam filtering tasks. This method takes the TF-IDF features of texts as input.
@@
  • Support Vector Machine (SVM) [19]: It is a popular classification algorithm that constructs an optimal hyperplane to separate different classes of samples. SVM maps data to a high-dimensional space and seeks a maximum-margin hyperplane to maximize the separation between classes. SVM performs well in handling high-dimensional data and non-linear classification problems, and different kernel functions can be used for different types of data. This method takes the TF-IDF features of texts as input.
@@
III-D Metrics
For classification models, the commonly used evaluation metrics are precision, recall, accuracy and F1 score [43]. Following the literature, notations $TP$, $TN$, $FP$, $FN$ represent the number of true positive, true negative, false positive and false negative, respectively. The accuracy (acc) refers to the proportion of correct predictions among the total number of predictions. It is calculated as follows:
$ACC=\frac{{TP+TN}}{{TN+TP+FN+FP}}$
Recall rate refers to the proportion of all true positive class (TP+FN) that is judged to be positive class (TP), that is, to judge whether a specific sample has been correctly classified, which is defined as follows:
$recall=\frac{{TP}}{{TN+TP}}$
The precision rate refers to the proportion of true positive classes (TP+FP) among all classes judged to be positive (TP+FP), namely, whether the class judged is completely correct.
$precision=\frac{{TP}}{{TP+FP}}$
The two indexes of Precision rate and Recall rate usually fluctuate, which is difficult to achieve both. They are mutually restricted in large-scale datasets, so comprehensive consideration is required. The most common method is F1-score, which is the weighted and average of precision and recall:
$F1=\frac{{2*P*R}}{{P+R}}$
@@ III-D Metrics
III-E Specific Settings
The experiment utilized the GPT3.5 API from Open AI to conduct experiments. We provide a prompt as input, excluding any historical dialogue or context, and instructing the model to disregard any previous context in the prompt [33]. We removed all line breaks and kept the message length under 100 words, in case the whole length was too long (GPT has a limit of 2000 tokens total). We specifically required the ChatGPT to output “ham” or “spam”, without any explanation.
@@ III-E Specific Settings
IV Results and Analysis
@@
TABLE II: A comparison of different methods for spam mail detection using macro-level precision, recall, F1 score and accuracy. The best and second-best results for each dataset and metrics are highlighted in bold and underlined, respectively.
@@
Precision Recall Accuracy F1 Precision Recall Accuracy F1
@@
Table II shows the performance of different models on both the English ESD and Chinese CSD datasets, in terms of macro-level precision, recall, F1 score and accuracy. From this table, BERT performs the best on the English ESD dataset, achieving the highest values on all four metrics, followed by DNN and LR. This result is pretty straightforward, as BERT is pre-trained and has the largest size of tunable parameters. In terms of macro F1 score, supervised learning methods (From naive Bayes to BERT) outperforms ChatGPT significantly, more than 10%. Based on the ESD dataset, the spam detection ability of ChatGPT is significantly weaker than supervised learning models, especially BERT. Fig. 1 illustrates the same information. This figure shows the performance of six methods on ESD in terms of four metrics (macro-level precision, recall, F1 score and accuracy). The ChatGPT with five-shot instances in the prompt is almost always below other lines.
@@
By comparing the performance of ChatGPT on ESD and CSD datasets, we find that it performs consistently on both datasets, with slightly higher F1 score on CSD.
IV-A Effect of Instance Number in the Prompt

### v2

@@
Abstract: Email continues to be a pivotal and extensively utilized communication medium within professional and commercial domains. Nonetheless, the prevalence of spam emails poses a significant challenge for users, disrupting their daily routines and diminishing productivity. Consequently, accurately identifying and filtering spam based on content has become crucial for cybersecurity. Recent advancements in natural language processing, particularly with large language models like ChatGPT, have shown remarkable performance in tasks such as question answering and text generation. However, its potential in spam identification remains underexplored. To fill in the gap, this study attempts to evaluate ChatGPT’s capabilities for spam identification in both English and Chinese email datasets. We employ ChatGPT for spam email detection using in-context learning, which requires a prompt instruction and a few demonstrations. We also investigate how the number of demonstrations in the prompt affects the performance of ChatGPT. For comparison, we also implement five popular benchmark methods, including naive Bayes, support vector machines (SVM), logistic regression (LR), feedforward dense neural networks (DNN), and BERT classifiers. Through extensive experiments, the performance of ChatGPT is significantly worse than deep supervised learning methods in the large English dataset, while it presents superior performance on the low-resourced Chinese dataset.
@@
Keywords: ChatGPT, Large Language Models, Spam email detection, Cybersecurity, In-context Learning
Mathematics Subject Classification: 03B65, 68T50, 91F20
1 Introduction
@@
Used as a swift and cost-efficient communication tool, email underpins global connectivity and is fundamental in various professional and personal activities, such as international electronic commerce (Hudak et al., 2017). As of 2020, the international email user base was approximately 4.037 billion. Due to its very nature of accessibility, email inevitably compromises on privacy (Ackerman et al., 1999), hence, exploited by corporations, organizations, and criminals alike for marketing and fraudulent purposes.
@@
Spam content predominantly encompasses deceptive information or advertisements (Jáñez-Martino et al., 2023), and occasionally includes harmful elements such as computer viruses, explicit material, or adverse information. The rampant spread of spam not only poses a nuisance to individuals but also instigates a squandering of network resources, in addition to eliciting a myriad of security threats (Ahmed et al., 2022). As such, the development of an efficacious model capable of accurately identifying and consequently filtering spam is of paramount importance (Samarthrao and Rohokale, 2022).
@@
Numerous methods have been established in the realm of spam detection and classification (Karim et al., 2019; Akinyelu, 2021). Conventional approaches encompass techniques such as character matching, the naive Bayes algorithm (Kumar, 2020), support vector machines (SVM) (Sharma and Kaur, 2016) and ensemble learning (Bhardwaj and Sharma, 2023). In the wake of deep learning advancements, a variety of deep models, including Convolutional Neural Networks (CNN), Recurrent Neural Networks (RNN), and BERT, have been employed in spam recognition (Seth and Biswas, 2017; Fazzolari et al., 2021; Manasa et al., 2024). For instance, Doshi et al. (2023) proposed a comprehensive dual-layer architecture for phishing and spam email detection. Zavrak and Yilmaz (2023) developed a hierarchical attention hybrid deep learning method for spam email detection.
@@
Recently large language models (LLMs) such as ChatGPT have shown impressive performance in many natural language processing tasks Radford et al. (2019), including question answering, text generation, etc. In-context learning Garg et al. (2022); Wei et al. (2022), central to these large models, refers to their prediction capabilities given a specific context; these models analyze and generate words or sentences based on the instruction (i.e., prompt), which aids their ability to coherently operate in tasks such as conversation, text completion, translation, and summarization. However, while ChatGPT has been used in various tasks of text processing, little research has been done on spam detection. To fill in the gap, we aim to explore how well ChatGPT performs at spam detection via in-context learning. We design experiments to systematically evaluate the performance of ChatGPT in both English and Chinese email datasets.
@@
2 Related Works
2.1 Spam Classification
@@
Based on the types of data used, existing spam classification algorithms can be broadly categorized into two groups: The first category involves recognition based on mail features. This approach involves extracting features from mail titles and employing supervised classifiers like the SVM model for classification (Khamis et al., 2020). Some studies have also considered the cumulative relative frequency of each spam feature to enhance the performance of supervised classifiers (Fazzolari et al., 2021). The feature-based spam recognition dataset is multidimensional, making it suitable for multiple classification models. Consequently, researchers have compared the performance of different machine learning models (Awad and ELseuofi, 2011), such as SVM, deep neural networks (DNN), and k-nearest neighbor (KNN) (Yeruva et al., 2022), in spam recognition tasks.
@@
The second category focuses on email content identification. For instance, the relaxed online SVM model can be utilized to effectively filter and classify mail content (Sculley and Wachman, 2007), naive Bayes classifiers have also demonstrated effectiveness in email classification (Ma et al., 2020). In recent years, deep learning algorithms have been increasingly employed in spam recognition. CNNs (Seth and Biswas, 2017), for example, have proven to be effective in identifying spam. The transfer learning of BERT models has also been employed for spam classification, yielding high accuracy (Fazzolari et al., 2021). Long Short-Term Memory (LSTM) models have shown better performance on short messages (SMS) datasets compared to machine learning classification algorithms like random forest and SVM (Raj et al., 2018).
2.2 GPT and Its Applications
@@
GPT (Radford et al., 2018) is one of the landmark models proposed on the basis of Transformer proposed by Vaswani in 2017 (Vaswani et al., 2017), another representative model is the BERT model proposed by (Devlin et al., 2018). On this basis, GPT2 (Radford et al., 2019) and GPT3 (Brown et al., 2020) have been proposed and widely used. Excellent performance was achieved tasks such as text generation, question and answer, translation, summary, and dialogue.
@@
ChatGPT (Ouyang et al., 2022) is an upgraded version of OpenAI in 2022 based on the GPT-3.5 model, which is optimized for conversation tasks, adding input and output of conversation history, and control of conversation policies. ChatGPT excels at conversational tasks and can have natural and fluid conversations with humans.
@@
Spam recognition is a text classification task. ChatGPT generally performs well on text classification tasks, and researchers have demonstrated that the model performs well on some text classification tasks (Liu et al., 2023a). GPT can be used to assist classification decision, by using chatgpt to extract structured and refined knowledge from knowledge graph as the characteristics of classification tasks. (Shi et al., 2023). It can also be used directly for text classification with appropriate Prompt design, such as for agricultural (Zhao et al., 2023), business news Lopez-Lira and Tang (2023) and even medical dialogues (Wang et al., 2023).
@@
A few research have explored the application of ChatGPT in text binary classification (Wang et al., 2023), which can provide some reference for spam recognition. In recent studies, ChatGPT was able to identify opinions, sentiment, and emotions contained in the text and categorize emotions. In addition, GPT also performed well in sentiment analysis of disease-related papers (Susnjak, 2023) and Twitter texts (Sharma et al., 2023), indicating that ChatGPT is capable of a wide range of sentiment analysis tasks for texts in different fields.
@@
Due to some uncertainty in the answers of large language models, (Borji, 2023), when using chatgpt for text classification tasks, it is necessary to carefully design hints and related strategies to make the answers as stable and effective as possible. If necessary, a small number of samples can be added to guide the model (Reiss, 2023), such as One-shot Prompting (Korini and Bizer, 2023),Few-shot Prompting (Zhao et al., 2021), and Chain-of-Thought (CoT) Prompting design (Liu et al., 2023b).
3 Spam Detection as an Optimization Problem
In this section, we formulate the spam email detection as a binary classification. Suppose the training dataset $\{(\textbf{x}_{1},y_{1}),(\textbf{x}_{2},y_{2}),\ldots,(\textbf{x}_{n},y_{n})\}$ consists of $n$ training examples, each of which has an email $(\textbf{x}_{i})$ and label $y_{i}\in\{0,1\}$. For a text encoder (feature extractor) $f_{\mathbf{\theta}}(\textbf{x})$, it extracts hidden representations from the input email x and it can be term frequency–inverse document frequency, BERT encoder (Devlin et al., 2018), feedforward neural networks (Manasa et al., 2024), and so on. The probability of the $i$-th email being spam is
$P(Y=1|\textbf{x}_{i})=\sigma(\textbf{w}^{T}f_{\mathbf{\theta}}(\textbf{x}_{i})+b)=\frac{\exp{(\textbf{w}^{T}f_{\mathbf{\theta}}(\textbf{x}_{i})+b)}}{1+\exp{(\textbf{w}^{T}f_{\mathbf{\theta}}(\textbf{x}_{i})+b)}},$ (3.1)
so we attempt to optimize the cross-entropy loss during the training stage:
$({\mathbf{\theta}}^{*},{\textbf{w}}^{*},{b}^{*})=\arg\min_{\mathbf{\theta},\textbf{w},b}\sum_{i=1}^{n}(y_{i}\log\sigma(\textbf{w}^{T}f_{\mathbf{\theta}}(\textbf{x}_{i})+b)+(1-y_{i})(1-\sigma(\textbf{w}^{T}f_{\mathbf{\theta}}(\textbf{x}_{i})+b))),$ (3.2)
where $\hat{\mathbf{\theta}},\hat{\textbf{w}},\hat{b}$ are the estimates of $\mathbf{\theta},\textbf{w}$ and $b$, respectively.
For the large language models (LLMs) like ChatGPT that has more than 100 Billion parameters, we do not need to tune the parameters, but to prompt them directly. The performance of a LLM depends on the design of prompts. We can tune the prompt on the validation dataset $S_{v}$,
$\textbf{p}^{*}=\arg\min_{\textbf{p}}\sum_{i\in{S}_{v}}I(\text{LLM}([\textbf{p},~{}\textbf{x}_{i}])\neq y_{i}),$ (3.3)
where $I(\cdot)$ is the indicator function, and it equals to 1 if the output of LLM is not $y_{i}$, otherwise it equals to 0. The optimization can be over the instruction of prompt or the training examples in the prompt.
4 Experimental Setup
@@ 4.1 Datasets
Table 1: Detail of Datasets
@@
The general information of the dataset is shown in Table 1. This table displays the size of two datasets, and the number of spam and ham emails in these datasets.
4.2 In-context Learning
In-context learning was popularized in the original GPT-3 paper (Brown et al., 2020) as a way to use language models to learn tasks given only a few examples in the instruction. During in-context learning, we give the language model a prompt that consists of a list of input-output pairs that demonstrate a task. At the end of the prompt, we append a test input and allow the language model to make a prediction just by conditioning on the prompt and predicting the next tokens.
To utilize ChatGPT via in-context learning, it is necessary to design an instruction, the so-called prompt.
4.2.1 Specification of Prompt
@@
Forget all your previous instructions, pretend you are a junk mail detection expert who try to identify whether a mail is spam mail. Answer “spam” if it is spam, “ham” if not. Now you can identify whether the mail is spam:
@@
4.2.2 Zero-shot Prompt
The output of ChatGPT is sometimes inconsistent, so we enhance the classification efficiency of the model by adjusting the design of prompt (Cao et al., 2023). On top of the observation from real-world questions, we propose an enhanced prompt template:
@@ 4.2.2 Zero-shot Prompt
We made this adjustment because ChatGPT sometimes had difficulty identifying email categories and gave responses like “neutral” or “ambiguous” under the zero-shot settings. Therefore, in order to achieve better binary classification, we utilize this prompt for the zero-shot experiments.
4.3 Baselines
@@
  • Naive Bayes: It is a classification algorithm based on Bayes’ theorem and feature independence assumption (Ma et al., 2020). It assumes that input features are independent of each other and uses prior probabilities of classes and conditional probabilities of features to calculate posterior probabilities for classification. Naive Bayes classifier is simple, computationally efficient, and widely used in text classification and spam filtering tasks. This method takes the TF-IDF features of texts as input.
@@
  • Support Vector Machine (SVM) (Khamis et al., 2020): It is a popular classification algorithm that constructs an optimal hyperplane to separate different classes of samples. SVM maps data to a high-dimensional space and seeks a maximum-margin hyperplane to maximize the separation between classes. SVM performs well in handling high-dimensional data and non-linear classification problems, and different kernel functions can be used for different types of data. This method takes the TF-IDF features of texts as input.
@@
4.4 Metrics
For classification models, the commonly used evaluation metrics are precision, recall, accuracy and F1 score (Pérez-Díaz et al., 2012). Following the literature, notations $TP$, $TN$, $FP$, $FN$ represent the number of true positive, true negative, false positive and false negative, respectively. The accuracy (Acc) refers to the proportion of correct predictions among the total number of predictions. It is calculated as follows:
$Acc=\frac{{TP+TN}}{{TN+TP+FN+FP}}.$
In the context of classification models, the recall (R) metric denotes the proportion of actual positive cases that are accurately identified by the model, which is defined as follows:
$R=\frac{{TP}}{{FN+TP}}.$
Conversely, the precision (P) metric represents the proportion of examples predicted as positive that are indeed truly positive, which is defined as follows,
$P=\frac{{TP}}{{TP+FP}}.$
Notably, the precision rate and recall rate often exhibit a trade-off phenomenon, where improvements in one metric may come at the expense of the other. This reciprocal relationship is particularly pronounced in large-scale datasets, necessitating a comprehensive consideration of both metrics. To strike a balance between these competing indices, the F1-score is commonly employed as a summary measure of model performance, which is defined as follows:
$F1=\frac{{2\times~{}P\times~{}R}}{{P+R}}.$
@@ 4.4 Metrics
4.5 Specific Settings
The experiment utilized the GPT3.5 API from OpenAI to conduct experiments. We provide a prompt as input, excluding any historical dialogue or context, and instructing the model to disregard any previous context in the prompt (Lopez-Lira and Tang, 2023). We removed all line breaks and kept the message length under 100 words, in case the whole length was too long (GPT has a limit of 2000 tokens total). We specifically required the ChatGPT to output “ham” or “spam”, without any explanation.
@@ 4.5 Specific Settings
5 Results and Analysis
5.1 Main Findings
@@
Table 2: A comparison of different methods for spam mail detection using macro-level precision (Prec.), recall (Rec.), F1 score and accuracy (Acc.). The best and second-best results for each dataset and metrics are highlighted in bold and underlined, respectively.
@@
Prec. Rec. Acc. F1 Prec. Rec. Acc. F1
@@
Table 2 shows the performance of different models on both the English ESD and Chinese CSD datasets, in terms of macro-level precision, recall, F1 score and accuracy. From this table, BERT performs the best on the English ESD dataset, achieving the highest values on all four metrics, followed by DNN and LR. This result is pretty straightforward, as BERT is pre-trained and has the largest size of tunable parameters. In terms of macro F1 score, supervised learning methods (From naive Bayes to BERT) outperforms ChatGPT significantly, more than 10%. Based on the ESD dataset, the spam detection ability of ChatGPT is significantly weaker than supervised learning models, especially BERT. Fig. 1 illustrates the same information. This figure shows the performance of six methods on ESD in terms of four metrics (macro-level precision, recall, F1 score and accuracy). The ChatGPT with five-shot instances in the prompt is almost always below other lines.
@@
By comparing the performance of ChatGPT on ESD and CSD datasets, we find that it performs consistently on both the English and Chinese datasets, ranging from 0.84 to 0.89 for accuracy and from 0.76 to 0.80 for F1 score.
5.2 Effect of Instance Number in the Prompt